OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9/5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

\[C=\frac{9}{5}(F-32)\] \[\frac{5}{9}C=F-32\] \[\frac{5}{9}C+32=F\]

OpenStudy (anonymous):

huh?

OpenStudy (ash2326):

For second part substitute C=4 in F we have \[F=\frac{5}{9}C+32\] C=4 \[F=\frac{5}{9}*4+32\] \[F=\frac{20}{9}+32\] we get \[F=2.22+32\] we get F=34.22 degree Fahrenheit